OpenStudy (anonymous):

"Sears Tower in Chicago is 1559 feet tall. What is its height in kilometers?" It's a three step dimensional analysis problem. ;p

OpenStudy (anonymous):

1 ft = 30.48 cm so 1559 ft = 1559 * 30.48 cm = 47518.32 cm 1 cm = 1/100000 km so 47518.32 cm = 47518.32 * (1/100000) km = 0.48 km approx.

OpenStudy (unklerhaukus):

1559 ft = 0.4752 km